OpenStudy (anonymous):

What are the formulas for momentum, force, velocity, distance, speed, time, weight, mass, acceleration??/

OpenStudy (anonymous):

note m = mass, p = momentum, a = acceleration, v = velocity, d = distance, a = acceleration, f=force, w=weight p = mv, f = ma v=d/t, d=vt t=d/v w = ma a = v/t